[Brahms-l] Shift report 20050129 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jan 29 2005 - 23:34:16 EST
1. RHIC was down for half of the shift, but when it came back up, the rates were nice.

2. Took some production data, but the MRS is still down.

3. No new problem were discovered.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 29 23:34:31 2005

This archive was generated by hypermail 2.1.8 : Sat Jan 29 2005 - 23:34:53 EST